OpenStudy (akohl):

@Abhisar A model rocket with a mass of 0.181 kg is launched into the air with an initial speed of 92 m/s. How much kinetic energy will the rocket have at a height of 164 m? Assume there is no wind resistance. 303 J 634 J 444 J 747 J

OpenStudy (akohl):

@Abhisar wasn't the Last question nuclear energy

OpenStudy (abhisar):

First use V^2=U^2+2as to find the final velocity. Then use 1/2mv^2 to find the kinetic energy.

OpenStudy (abhisar):

Yes, answer to last question was nuclear energy.

OpenStudy (akohl):

thanks a is gravity right

OpenStudy (abhisar):

a=9.8

OpenStudy (akohl):

I got 1057

OpenStudy (akohl):

But that's not an option

OpenStudy (abhisar):

What is the value of velocity u have got ?

OpenStudy (akohl):

108.1

OpenStudy (abhisar):

v^2=(92)^2 - 2*9.8*164 -->v^2=8464-3214.4 -->v^2=5249.6

OpenStudy (abhisar):

I think the answer should be 3rd option !

OpenStudy (akohl):

K I used that and got 475j not 444

OpenStudy (abhisar):

yes !

OpenStudy (akohl):

Maybe the computer screwed up the question thanks

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)

OpenStudy (abhisar):

u sure u r not missing any fact from the question ?

OpenStudy (akohl):

No I checked